Mary is half as old as Tom. George is 3 years older than tom. The sum of Mary,tom,and george's ages is 58..How old is tom?
Answer by Maths68(1474)

Let
Present age of George = g
Present age of Tom =t
Present age of Mary = m
Mary is half as old as Tom
m=1/2*t
m=t/2...............(1)
George is 3 years older than tom
g=t+3...............(2)
The sum of Mary,tom,and george's ages is 58
g+t+m=58
Put the value of g and m in above euqation
(t+3)+t+(t/2)=58
t+3+t+(t/2)=58
Multiply by 2 both sides
2*(t+3+t+(t/2))=2*58
2t+6+2t+t=116
5t=116-6
5t=110
5t/5=110/5
t=22


Tom is 22 years old
For George and Mary's age put the value of t in (1) and (2)
m=t/2=22/2=11
g=22+3=25
George is 25 and Mary is 11 years old